you are in charge of building the roof for a warehouse. it is a requirement that the pitch or slope of the roof rises at least 3 feet for every 15 feet of horizontal distance.
a. what is the required pitch of slope of the roof?
b. if the warehouse is 120 feet wide, what is the minimum required height of the roof?
c. use the pythagorean theorem to determine the lenght of one side of the roof.
d. explain how a roof height of 15 feet meets the required pitch for the roof.

Answer by stanbon(75887)

it is a requirement that the pitch or slope of the roof rises at least 3 feet for every 15 feet of horizontal distance.
a. what is the required pitch of slope of the roof?
pitch = 3/15 = 1/5
--------------------------------
b. if the warehouse is 120 feet wide, what is the minimum required height of the roof?
x/120 = 1/3
x = 120/3
x= 40 ft (this is the minimum height)
------------------------------
c. use the pythagorean theorem to determine the length of one side of the roof.
120^2 = 40^2 + x^2
x^2 = 12000
x = 113.137 ft (the length of one side of the roof)
--------------------------------
d. explain how a roof height of 15 feet meets the required pitch for the roof.
It does if the the run of the roof is 45 ft.
===============================
Cheers,
stan H.